OpenStudy (anonymous):

From a barrel of colored marbles, you randomly select 7 blue, 5 yellow, 8 red, 4 green, and 6 purple marbles. A: Find the probability of randomly selecting a marble that is not yellow. Write your answer insimplest form. ~A:1/30B:5/6 C:2/15 D:1/6 B: Find the expiramental probability of randomly selecting a marble that is either green or purple. Write your answer in simplest form. ~A:1/10 ~B:1/5 ~C:1/3 ~ D:2/15 Will become Fan, give medal, and write testimony:D

OpenStudy (anonymous):

Its a little diorganized, sorry:(

OpenStudy (anonymous):

7+5+8+4+6=30 total marbles 5 yellows 30-5=25 marbles that aren't yellow. 25/30 marbles aren't yellow, so simplified, the answer is B. 5/6.

OpenStudy (anonymous):

Again, there are 30 marbles. 4+6=10 green or purple marbles. 10/30=1/3.

OpenStudy (anonymous):

alright Thanks :)